# Partial information decomposition as information bottleneck

### v1

@@ III.4 Examples
Figure 2(d) shows the RB curves for the different sources. Source $X_{4}$ does not contribute to prediction at any level of compression, in accord with the fact that $I(Q;Y|S=s)\leq I(X_{4};Y)=0$. Sources $X_{1}$ and $X_{2}$ provide the same amount of prediction and compression at all points, up to the maximum $I(X_{2};Y)\approx 0.531$. Source $X_{3}$ provides the same amount of prediction and compression as sources $X_{1}$ and $X_{2}$, until it hits its maximum $I(X_{3};Y)\approx 0.278$. At this point, $X_{3}$ splits off from sources $X_{1}$ and $X_{2}$ and its compression contribution decreases to 0; this is compensated by the increasing compression cost of sources $X_{1}$ and $X_{2}$. Interestingly, we see that the RB curves of the individual sources may not be concave, unlike the overall RB curve.
Observe that the RB curves for the individual sources allow us to identify that sources $X_{1}$ and $X_{2}$ provide the most redundant information.
Figure 3: RB analysis for the system with a 3-spin target (Example 4). (a) and (b): Prediction and compression values found by optimizing the RB Lagrangian (14) at different $\beta$. (c) The RB curve shows the tradeoff between optimal compression and prediction values; marker colors correspond to $\beta$ values, as in (a) and (b). All possible bottleneck variables $Q$ fall within the accessible grey region. (d) The RB curve is decomposed into prediction and compression contributions from different sources. Line colors and styles indicate different sources.
Example 4.
@@
In our final example, the target consists of 3 binary spins with a uniform distribution, so $Y=(Y_{1},Y_{2},Y_{3})$ and $p_{Y}(y)=1/8$ for all $y$. There are three sources, each of which contains two binary spins. Sources $X_{1}$ and $X_{2}$ are both equal to the first two spins of the target $Y$, $X_{1}=X_{2}=(Y_{1},Y_{2})$. Source $X_{3}$ is equal to the first and last spin of the target, $X_{3}=(Y_{1},Y_{3})$. Each source provides $I(Y;X_{s})=2\text{ bits}$ of mutual information about the target.
@@
The Blackwell redundancy $I_{\cap}$ is 1 bit, reflecting the fact that there is a single binary spin that is included in all sources ($Y_{1}$).
@@
We perform our RB analysis and plot RB prediction values in Figure 3(a) and compression values in Figure 3(b), as found by optimizing the RB Lagrangian at different $\beta$. At small $\beta$, prediction converges to Blackwell redundancy, $I(Q;Y|S)=I_{\cap}=1$, and $I(Q;S|Y)=0$. At large $\beta$, the prediction is equal to the maximum $I(Z;Y|S)=2\text{ bit}$, and compression is equal to $I(Z;S|Y)\approx 0.459$. As in the previous example, the RB curve undergoes phase transitions as different optimal strategies are uncovered at different $\beta$. Figure 3(c) shows the RB curve.
@@ III.5 Continuity
In the final section of our theoretical analysis, we consider the continuity of RB as a function of the underlying probability distribution.
It is known Blackwell redundancy $I_{\cap}$ can be discontinuous as a function of the probability distribution of the target and source channels [4]. In Ref. [4], we explain the origin of this discontinuity in geometric terms and provide sufficient conditions for Blackwell redundancy to be continuous. Nonetheless, the discontinuity of $I_{\cap}$ is sometimes seen as an undesired property.
On the other hand, RB is continuous for all $R>0$.
@@
For finite dimensional systems and $R>0$, $I_{\text{RB}}(R)$ is a continuous function of the probability values $p_{X_{s}|Y}(x|y)$, $p_{Y}(y)$, and $\nu_{S}(s)$.
@@
Example 5.
We consider the COPY gate, a standard example in the PID literature. Here there are two binary-valued sources jointly distributed according to
@@ IV Iterative algorithm
We now provide an iterative algorithm to solve the RB optimization problem. This algorithm is conceptually similar to the Blahut-Arimoto algorithm, originally employed for rate-distortion problems and later adapted to solve the original IB problem [6]. A Python implementation of our algorithm is available at https://github.com/artemyk/pid-as-ib; there we also provide updated code to compute Blackwell redundancy (applicable to small systems).
To begin, we consider the RB Lagrangian optimization problem, Eq. (14). We rewrite this optimization problem using KL divergences,
$F_{\text{RB}}(\beta)=\sup_{r_{Q|SZ}}D(r_{Y|QS}\|p_{Y|S})-\frac{1}{\beta}D(r_{Q|SY}\|r_{Q|Y}).$ (23)
Here, notation like $r_{Q|Y}$, $r_{Y|QS}$, etc. refers to distributions that include $Q$ and therefore depend on the optimization variable $r_{Q|SZ}$, while notation like $p_{Y|S}$ refers to distributions that do not depend on $Q$ and are not varied under the optimization. Every choice of conditional distribution $r_{Q|SZ}$ induces a joint distribution $r_{YSZQ}=p_{YSZ}r_{Q|SZ}$ via Eq. (10).
@@
$D(r_{Y|QS}\|p_{Y|S})$
$\qquad=D(r_{Y|QS}\|p_{Y|S})-\min_{\omega_{YSZQ}}D(r_{Y|QS}\|\omega_{Y|QS})$
$\qquad=\max_{\omega_{YSZQ}}\mathbb{E}_{p_{YSZ}r_{Q|SZ}}\left[\ln\frac{\omega(y|q,s)}{p(y|s)}\right].$
where $\mathbb{E}$ indicates expectation and we introduced the variational distribution $\omega_{YSZQ}$. The maximum is achieved by $\omega_{YSZQ}=r_{YSZQ}$, which gives $\omega_{Y|QS}=r_{Y|QS}$. We rewrite the second KL term in Eq. (23) as
@@
As before, we introduced the variational distribution $\omega_{YSZQ}$, where the minimum is again achieved by $\omega_{YSZQ}=r_{YSZQ}$. The term $r_{Q|SY}r_{Z|SYQ}$ can be rewritten as
$r(q|s,y)r(z|s,y,q)$ $=\frac{r(z,s,y,q)}{p(s,y)}$
$=\frac{r(q|s,z)p(z,s,y)}{p(s,y)}=r(q|s,z)p(z|s,y)$
@@
$F_{\text{RB}}(\beta)=\sup_{r_{Q|ZS},\omega_{YSZQ}}\mathbb{E}_{p_{YSZ}r_{Q|SZ}}\left[\ln\frac{\omega(y|q,s)}{p(y|s)}\right]\\ -\frac{1}{\beta}D(r_{Q|SZ}p_{Z|SY}\|\omega_{Q|Y}\omega_{Z|SYQ}).$ (24)
We can now optimize this objective in an iterative and alternating manner with respect to $r_{Q|SZ}$ and $\omega_{YSZQ}$. Formally, let $\mathcal{L}(r_{Q|SZ},\omega_{YSZQ})$ refer to the objective in Eq. (24). Then, starting from some initial guess $r_{Q|SZ}^{(0)}$, we generate a sequence of solutions
@@
Each optimization problem can be solved in closed form. As already mentioned, $\omega_{YSZQ}^{(t+1)}=r_{YSZQ}^{(t)}=r_{Q|SZ}^{(t)}p_{SZY}$ in (25). The optimization (26) can be solved by taking derivatives and setting to 0, giving
$r^{(t+1)}(q|s,z)\propto\\ e^{\sum_{y}p(y|s,z)\Big{[}\beta\ln\omega^{(t)}(y|q,s)-\ln\frac{p(z|s,y)}{\omega^{(t)}(q|y)\omega^{(t)}(z|s,y,q)}\Big{]}},$
@@
Each iteration increases the value of objective $\mathcal{L}$. Since the objective is upper bounded by $I(Z,Y|S)$, the algorithm is guaranteed to converge. However, as in the case of the original IB problem, the objective is not convex jointly in both arguments, so the algorithm may converge to a local maximum or a saddle point, rather than a global maximum. This can be partially alleviated by running the algorithm several times starting from different initial guesses $r_{Q|SZ}^{(0)}$.
For non-strictly-concave RB curve, it is more appropriate to optimize the exponential RB Lagrangian 15 or another objective that combines prediction and compression terms in a nonlinear manner [18, 19]. The algorithm described above can be used with such objectives after a slight modification. For instance, for the exponential RB Lagrangian, we may rewrite the optimization using an additional variational distribution $\omega_{YSZQ}$ as
$F_{\text{RB}}^{\exp}(\beta)=\sup_{r_{Q|SZ},\omega_{YSZQ}}\mathbb{E}_{p_{YSZ}r_{Q|SZ}}\left[\ln\frac{\omega(y|q,s)}{p(y|s)}\right]\\ -\frac{1}{\beta}e^{D(r_{Q|ZS}p_{Z|SY}\|\omega_{Q|Y}\omega_{Z|SYQ})}.$ (27)
@@
where $\beta^{(t)}=\beta e^{-I_{r^{(t)}}(Q;S|Y)}$. (Observe that, unlike the squared Lagrangian [18], the exponential Lagrangian leads to an effective inverse temperature $\beta^{(t)}$ that is always finite and converges to $\beta$ for $I_{r^{(t)}}(Q;S|Y)\to 0$).
When computing an entire RB curve, as in Figure 1(a)-(c), we found good results by re-using the optimal $r_{Q|SZ}$ found for one $\beta$ as the initial guess at higher $\beta$, which may be understood as a kind of annealing. In addition, for quantifying the value of the RB function $I_{\text{RB}}(R)$ at a fixed $R$, as in Figure 4, we approximated $I_{\text{RB}}(R)$ via a linear interpolation of the RB prediction and compression values recovered from the RB Lagrangian at varying $\beta$.
@@ V Discussion
In this paper, we proposed a generalization of Blackwell redundancy, termed the redundancy bottleneck (RB), formulated as an information-bottleneck-type tradeoff between prediction and compression. We study some implications of this formulation and propose an efficient numerical algorithm to solve the RB optimization problem.
We briefly mention some interesting directions for future work.
The first direction concerns our iterative algorithm. The algorithm is only applicable to systems where it is possible to enumerate the outcomes of the joint distribution $p_{QYSZ}$. This is impractical for discrete-valued variables with very many outcomes and for continuous-valued variables, e.g., as commonly found in machine learning settings where the data is a set of high-dimensional samples. In future work, it would be useful to develop RB algorithms suitable for such datasets, possibly by exploiting the kinds of variational techniques that have recently gained traction in machine learning applications of IB [11, 12, 13].
The second direction would explore connections between the RB and other information-theoretic objectives for representation learning. To our knowledge, the RB problem is novel to the literature; however, it has some similarities to objectives such as conditional entropy bottleneck [13], multi-view IB [22], and the privacy funnel and its variants [50], among others. Showing formal connections between these objectives would be of theoretical and practical interest, and could lead to new interpretations of the concept of PID redundancy. Another direction would explore the relationship between RB and information-theoretic measures of causality [51, 52]. Specifically, if the source identity $S$ is used to represent a confounding nuisance variable, then redundancy could provide a measure of causal information flow that is invariant to the choice of the nuisance.
Finally, one of the central ideas of this paper is to treat the source identity as a random variable in its own right, allowing us to consider what information different bottleneck variables reveal about the source. In this way, we convert the search for topological or combinatorial structure in multivariate systems into an interpretable and differential information-theoretic objective. This general idea may be useful in other problems that concern how information is distributed in complex systems. Possible applications include the study of other PID measures such as synergy [4], information-theoretic measures of modularity [53, 54], and measures of higher-order dependency [55, 56].
@@
We provide proofs of the theorems in the main text.
@@
Let $V_{1}$ indicate the supremum in Eq. (7) and $V_{2}$ the supremum in the optimization of Theorem 1, given some $\nu_{S}(s)$ with full support. We prove that $V_{1}=V_{2}$.
We will use the fact that, for any distribution that has the form of Eq. (10) and obeys $I(Q;S|Y)=0$, the following holds:
@@
Here we used the Markov condition $Q-Y-S$ as well as $I(Q;S)=H(Q)-H(Q|S)=0$, as follows from Eq. (9) and the data processing inequality.
Let $Q$ be a feasible random variable that comes within $\epsilon\geq 0$ of the objective in (7), $I(Q;Y)\geq V_{1}-\epsilon$. Define the joint distribution
@@
whenever $z\in\mathcal{X}_{s}$ (otherwise $p_{QYSZ}(q,y,s,z)=0$). Here, we used the channels $\kappa_{Q|X_{s}}$ associated with the Blackwell relation $Q\preceq_{Y}X_{s}$, so that $p_{Q|Y}=\kappa_{Q|X_{s}}\circ p_{X_{s}|Y}$. Under the distribution $p_{QYSZ}$, the Markov conditions $Q-(S,Z)-Y$ and $Q-Y-S$ hold, the latter since
@@
Therefore, this distribution has the form of Eqs. (10) and (8) and satisfies the constraints in Theorem 1. Using Eq. (28), we then have
@@
Conversely, let $p_{YSZQ}$ be a feasible joint distribution for the optimization of Theorem 1 that comes within $\epsilon\geq 0$ of the supremum, $I(Q;Y|S)\geq V_{2}-\epsilon$. Using the form of this joint distribution from Eq. (10), we can write
@@
In $(a)$ we used $I(Q;S|Y)=0$, in $(b)$ we used $Q-(S,Z)-Y$, and in $(c)$ we used that $p_{Z|YS=s}=p_{X_{s}|Y}$. This implies that $p_{Q|Y}\preceq p_{X_{s}|Y}$ for all $s$. Therefore $p_{Q|Y}$ satisfies the constraints in Eq. (7), so $I(Q;Y)\leq V_{1}$. Combining with Eq. (28) implies
@@
To show concavity, assume that the supremum is achieved in the optimization problem (11) (if it is not, the same argument follow by using an additional limiting argument). Then, consider any two optimal points on the RB curve, $(R,I_{\text{RB}}(R))$ and $(R^{\prime},I_{\text{RB}}(R^{\prime}))$, with corresponding optimal variables $Q$ and $Q^{\prime}$. Without loss of generality, suppose that both outcome variables have the same set of outcomes $\mathcal{Q}$. Then, we define a new random variable $Q_{\lambda}$ with outcomes $\mathcal{Q}_{\lambda}=\{1,2\}\times\mathcal{Q}$, as well as a family of conditional distributions parameterized by $\lambda\in[0,1]$,
@@
Basically, we define $Q_{\lambda}$ via a disjoint convex mixture of $Q$ and $Q^{\prime}$ onto non-overlapping subspaces, with $\lambda$ being the mixing parameter. With a bit of algebra, it can be verified that for every $\lambda$,
@@
$=\lambda R+(1-\lambda)R^{\prime}$
@@
$=\lambda I_{\text{RB}}(R)+(1-\lambda)I_{\text{RB}}(R^{\prime})$
Since $I_{\cap}^{R}$ is defined via a maximization, we have
$I_{\text{RB}}(\lambda R+(1-\lambda)R^{\prime})$ $\geq I(Q_{\lambda};Y)$
$=\lambda I_{\text{RB}}(R)+(1-\lambda)I_{\text{RB}}(R^{\prime}),$
which proves concavity. ∎

### v2

@@ III.4 Examples
As expected, source $X_{4}$ does not contribute to the prediction at any level of compression, in accord with the fact that $I(Q;Y|S=s)\leq I(X_{4};Y)=0$. Sources $X_{1}$ and $X_{2}$ provide the same amount of prediction and compression at all points, up to the maximum $I(X_{1};Y)=I(X_{2};Y)\approx 0.531$. Source $X_{3}$ provides the same amount of prediction and compression as sources $X_{1}$ and $X_{2}$, until it hits its maximum prediction $I(X_{3};Y)\approx 0.278$. As shown in Figure 2(d), at this point, $X_{3}$ splits off from sources $X_{1}$ and $X_{2}$ and its compression contribution decreases to 0; this is compensated by increasing the compression cost of sources $X_{1}$ and $X_{2}$. The same behavior can also be seen in Figure 2(a) and (b), where we see that the solutions undergo phase transitions as different optimal strategies are uncovered at increasing $\beta$. Importantly, by considering the prediction/compression contributions from the the individual sources, we can identify that sources $X_{1}$ and $X_{2}$ provide the most redundant information.
Let us comment on the somewhat surprising fact that, at larger $\beta$, the compression cost of $X_{3}$ decreases—even while its prediction contribution remains constant and the prediction contribution from $X_{1}$ and $X_{2}$ increases. At first glance, this appears counter-intuitive if one assumes that, in order to increase prediction from $X_{1}$ and $X_{2}$, the bottleneck channel $p_{Q|Y}$ should approach $p_{X_{1}|Y}=p_{X_{2}|Y}$, thereby increasing the deficiency $\delta_{D}(p_{X_{3}|Y},p_{Q|Y})$ and the compression cost of $X_{3}$ via the bound (23). In fact, this is not the case, because the prediction is quantified via the conditional mutual information $I(Q;Y|S)$, not the mutual information $I(Q;Y)$. Thus, it is possible that the prediction contributions from $X_{1}$ and $X_{2}$ are large, even when the bottleneck channel $p_{Q|Y}$ does not closely resemble $p_{X_{1}|Y}=p_{X_{2}|Y}$.
More generally, this example shows that it is possible for the prediction contribution from a given source to stay the same, or even increase, while its compression cost decreases. In other words, as can be seen from Figure 2(d), it is possible for the RB curves of the individual sources to be non-concave and non-monotonic. It is only the overall RB curve, Figure 2(c), representing the optimal prediction–compression tradeoff on average, that must be concave and monotonic.
@@
Example 4. In our final example, the target consists of three binary spins with a uniform distribution, so $Y=(Y_{1},Y_{2},Y_{3})$ and $p_{Y}(y)=1/8$ for all $y$. There are three sources, each of which contains two binary spins. Sources $X_{1}$ and $X_{2}$ are both equal to the first two spins of the target $Y$, $X_{1}=X_{2}=(Y_{1},Y_{2})$. Source $X_{3}$ is equal to the first and last spin of the target, $X_{3}=(Y_{1},Y_{3})$.
@@
Each source provides $I(Y;X_{s})=2\text{ bits}$ of mutual information about the target. The Blackwell redundancy $I_{\cap}$ is 1 bit, reflecting the fact that there is a single binary spin that is included in all sources ($Y_{1}$).
@@
We performed our RB analysis and plot the RB prediction values in Figure 3(a) and the compression values in Figure 3(b), as found by optimizing the RB Lagrangian at different $\beta$. At small $\beta$, the prediction converges to the Blackwell redundancy, $I(Q;Y|S)=I_{\cap}=1$, and $I(Q;S|Y)=0$. At large $\beta$, the prediction is equal to the maximum $I(Z;Y|S)=2\text{ bit}$, and compression is equal to $I(Z;S|Y)\approx 0.459$. Figure 3(c) shows the RB curve. As in the previous example, the RB curve undergoes phase transitions as different optimal strategies are uncovered at different $\beta$.
In Figure 3(a) and (b), we show the additive contributions to the prediction and compression terms from the individual sources, $\nu_{S}(s)I(Q;Y|S=s)$ and $\nu_{S}(s)I(Q;S=s|Y)$, as shaded regions. We also show the resulting RB curves for individual sources in Figure 3(d).
@@ III.5 Continuity
It is known that the Blackwell redundancy $I_{\cap}$ can be discontinuous as a function of the probability distribution of the target and source channels [4]. In Ref. [4], we explain the origin of this discontinuity in geometric terms and provide sufficient conditions for Blackwell redundancy to be continuous. Nonetheless, the discontinuity of $I_{\cap}$ is sometimes seen as an undesired property.
On the other hand, as we show in this section, the value of RB is continuous in the probability distribution for all $R>0$.
@@
For finite-dimensional systems and $R>0$, $I_{\text{RB}}(R)$ is a continuous function of the probability values $p_{X_{s}|Y}(x|y)$, $p_{Y}(y)$, and $\nu_{S}(s)$.
@@
Example. We consider the COPY gate, a standard example in the PID literature. Here, there are two binary-valued sources jointly distributed according to
@@ IV Iterative Algorithm
We provide an iterative algorithm to solve the RB optimization problem. This algorithm is conceptually similar to the Blahut–Arimoto algorithm, originally employed for rate distortion problems and later adapted to solve the original IB problem [6]. A Python implementation of our algorithm is available at https://github.com/artemyk/pid-as-ib; there, we also provide updated code to exactly compute Blackwell redundancy (applicable to small systems).
To begin, we consider the RB Lagrangian optimization problem, Eq. (15). We rewrite this optimization problem using the KL divergence:
$F_{\text{RB}}(\beta)=\max_{r_{Q|SZ}}D(r_{Y|QS}\|p_{Y|S})-\frac{1}{\beta}D(r_{Q|SY}\|r_{Q|Y}).$ (25)
Here, notation like $r_{Q|Y}$, $r_{Y|QS}$, etc., refers to distributions that include $Q$ and therefore depend on the optimization variable $r_{Q|SZ}$, while notation like $p_{Y|S}$ refers to distributions that do not depend on $Q$ and are not varied under the optimization. Every choice of conditional distribution $r_{Q|SZ}$ induces a joint distribution $r_{YSZQ}=p_{YSZ}r_{Q|SZ}$ via Eq. (10).
@@
$D(r_{Y|QS}\|p_{Y|S})$ $=D(r_{Y|QS}\|p_{Y|S})-\min_{\omega_{YSZQ}}D(r_{Y|QS}\|\omega_{Y|QS})$
$=\max_{\omega_{YSZQ}}\mathbb{E}_{p_{YSZ}r_{Q|SZ}}\left[\ln\frac{\omega(y|q,s)}{p(y|s)}\right].$
where $\mathbb{E}$ indicates the expectation, and we introduced the variational distribution $\omega_{YSZQ}$. The maximum is achieved by $\omega_{YSZQ}=r_{YSZQ}$, which gives $\omega_{Y|QS}=r_{Y|QS}$. We rewrite the second KL term in Eq. (25) as
@@
Here, we introduce the variational distribution $\omega_{YSZQ}$, where the minimum is achieved by $\omega_{YSZQ}=r_{YSZQ}$. The term $r_{Q|SY}r_{Z|SYQ}$ can be rewritten as
$r(q|s,y)r(z|s,y,q)=\frac{r(z,s,y,q)}{p(s,y)}\\ =\frac{r(q|s,z)p(z,s,y)}{p(s,y)}=r(q|s,z)p(z|s,y)$
@@
$F_{\text{RB}}(\beta)=\max_{r_{Q|ZS},\omega_{YSZQ}}\mathbb{E}_{p_{YSZ}r_{Q|SZ}}\left[\ln\frac{\omega(y|q,s)}{p(y|s)}\right]\\ -\frac{1}{\beta}D(r_{Q|SZ}p_{Z|SY}\|\omega_{Q|Y}\omega_{Z|SYQ}).$ (26)
We now optimize this objective in an iterative and alternating manner with respect to $r_{Q|SZ}$ and $\omega_{YSZQ}$. Formally, let $\mathcal{L}(r_{Q|SZ},\omega_{YSZQ})$ refer to the objective in Eq. (26). Then, starting from some initial guess $r_{Q|SZ}^{(0)}$, we generate a sequence of solutions
@@
Each optimization problem can be solved in closed form. As already mentioned, the optimizer in Eq. (27) is
$\omega_{YSZQ}^{(t+1)}=r_{YSZQ}^{(t)}=r_{Q|SZ}^{(t)}p_{SZY}.$
The optimization (28) can be solved by taking derivatives, giving
$r^{(t+1)}(q|s,z)\propto e^{\sum_{y}p(y|s,z)\Big{[}\beta\ln\omega^{(t)}(y|q,s)-\ln\frac{p(z|s,y)}{\omega^{(t)}(q|y)\omega^{(t)}(z|s,y,q)}\Big{]}}$
@@
Each iteration increases the value of objective $\mathcal{L}$. Since the objective is upper bounded by $I(Z;Y|S)$, the algorithm is guaranteed to converge. However, as in the case of the original IB problem, the objective is not jointly convex in both arguments, so the algorithm may converge to a local maximum or a saddle point, rather than a global maximum. This can be partially alleviated by running the algorithm several times starting from different initial guesses $r_{Q|SZ}^{(0)}$.
When the RB is not strictly concave, it is more appropriate to optimize the exponential RB Lagrangian (16) or another objective that combines the prediction and compression terms in a nonlinear manner [18, 19]. The algorithm described above can be used with such objectives after a slight modification. For instance, for the exponential RB Lagrangian, we modify (26) as
$F_{\text{RB}}^{\exp}(\beta)=\max_{r_{Q|SZ},\omega_{YSZQ}}\mathbb{E}_{p_{YSZ}r_{Q|SZ}}\left[\ln\frac{\omega(y|q,s)}{p(y|s)}\right]\\ -\frac{1}{\beta}e^{D(r_{Q|ZS}p_{Z|SY}\|\omega_{Q|Y}\omega_{Z|SYQ})}.$ (29)
@@
where $\beta^{(t)}=\beta e^{-I_{r^{(t)}}(Q;S|Y)}$ is an effective inverse temperature. (Observe that, unlike the squared Lagrangian [18], the exponential Lagrangian leads to an effective inverse temperature $\beta^{(t)}$ that is always finite and converges to $\beta$ as $I_{r^{(t)}}(Q;S|Y)\to 0$.)
When computing an entire RB curve, as in Figure 1(a)-(c), we found good results by annealing, that is by re-using the optimal $r_{Q|SZ}$ found for one $\beta$ as the initial guess at higher $\beta$. For quantifying the value of the RB function $I_{\text{RB}}(R)$ at a fixed $R$, as in Figure 4, we approximated $I_{\text{RB}}(R)$ via a linear interpolation of the RB prediction and compression values recovered from the RB Lagrangian at varying $\beta$.
@@ V Discussion
In this paper, we propose a generalization of Blackwell redundancy, termed the redundancy bottleneck (RB), formulated as an information-bottleneck-type tradeoff between prediction and compression. We studied some implications of this formulation and proposed an efficient numerical algorithm to solve the RB optimization problem.
We briefly mention some directions for future work.
The first direction concerns our iterative algorithm. The algorithm is only applicable to systems where it is possible to enumerate the outcomes of the joint distribution $p_{QYSZ}$. This is impractical for discrete-valued variables with very many outcomes, as well as continuous-valued variables as commonly found in statistical and machine learning settings. In future work, it would be useful to develop RB algorithms suitable for such datasets, possibly by exploiting the kinds of variational techniques that have recently gained traction in machine learning applications of IB [11, 12, 13].
The second direction would explore connections between the RB and other information-theoretic objectives for representation learning. To our knowledge, the RB problem is novel to the literature. However, it has some similarities to existing objectives, including among others the conditional entropy bottleneck [13], multi-view IB [22], and the privacy funnel and its variants [49]. Showing formal connections between these objectives would be of theoretical and practical interest, and could lead to new interpretations of the concept of PID redundancy.
Another direction would explore the relationship between RB and information-theoretic measures of causality [50, 51]. In particular, if the different sources represent some mutually exclusive conditions—such as the age group example provided at the end of Section II—then redundancy could serve as a measure of causal information flow that is invariant to the conditioning variable.
Finally, one of the central ideas of this paper is to treat the identity of the source as a random variable in its own right, which allows us to consider what information different bottleneck variables reveal about the source. In this way, we convert the search for topological or combinatorial structure in multivariate systems into an interpretable and differential information-theoretic objective. This technique may be useful in other problems that consider how information is distributed among variables in complex systems, including other PID measures such as synergy [4], information-theoretic measures of modularity [52, 53], and measures of higher-order dependency [54, 55].
Acknowledgements.
I thank Nihat Ay, Daniel Polani, Fernando Rosas, and especially, André Gomes for useful feedback. I also thank the organizers of the “Decomposing Multivariate Information in Complex Systems” (DeMICS 23) workshop at the Max Planck Institute for the Physics of Complex Systems (Dresden, Germany), which inspired this work. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 101068029.
@@ Acknowledgements.
We provide proofs of the theorems in the main text. Throughout, we use $D$ for the Kullback–Leibler (KL) divergence, $H$ for the Shannon entropy, and $I$ for the mutual information.
@@
We begin by proving a slightly generalized version of Theorem 1, that is we show the equivalence between the two optimization problems:
$I_{\cap}$ $=\sup_{Q}\,I(Q;Y)\quad\textrm{where}\quad Q\preceq_{Y}X_{s}\;\forall s$ (30)
$I_{\cap}$ $=\sup_{Q:Q-(Z,S)-Y}I(Q;Y|S)\quad\textrm{where}\quad I(Q;S|Y)=0.$ (31)
The slight generalization comes from replacing $\max$ by $\sup$, so that the result also holds for systems with infinite-dimensional sources, where the supremum is not guaranteed to be achieved. For finite-dimensional systems, the supremum is always achieved, and we reduce to the simpler case of Eqs. (7) and (11).
@@
Let $V_{1}$ indicate the supremum in Eq. (30) and $V_{2}$ the supremum in Eq. (31), given some $\nu_{S}(s)$ with full support. We prove that $V_{1}=V_{2}$.
We will use that, for any distribution that has the form of Eq. (10) and obeys $I(Q;S|Y)=0$, the following holds:
@@
Here, we used the Markov condition $Q-Y-S$, as well as $I(Q;S)=H(Q)-H(Q|S)=0$, as follows from Eq. (9) and the data-processing inequality.
Let $Q$ be a feasible random variable that comes within $\epsilon\geq 0$ of the objective in (30), $I(Q;Y)\geq V_{1}-\epsilon$. Define the joint distribution:
@@
whenever $z\in\mathcal{X}_{s}$, otherwise $p_{QYSZ}(q,y,s,z)=0$. Here, we used the channels $\kappa_{Q|X_{s}}$ associated with the Blackwell relation $Q\preceq_{Y}X_{s}$, so that $p_{Q|Y}=\kappa_{Q|X_{s}}\circ p_{X_{s}|Y}$. Under the distribution $p_{QYSZ}$, the Markov conditions $Q-(S,Z)-Y$ and $Q-Y-S$ hold, the latter since
@@
Therefore, this distribution has the form of Eqs. (10) and (8) and satisfies the constraints in Eq. (31). Using Eq. (32), we then have
@@
Conversely, let $p_{YSZQ}$ be a feasible joint distribution for the optimization of Eq. (31) that comes within $\epsilon\geq 0$ of the supremum, $I(Q;Y|S)\geq V_{2}-\epsilon$. Using the form of this joint distribution from Eq. (10), we can write
@@
In $(a)$, we used $I(Q;S|Y)=0$; in $(b)$, we used $Q-(S,Z)-Y$; in $(c)$, we used that $p_{Z|YS=s}=p_{X_{s}|Y}$. This implies that $p_{Q|Y}\preceq p_{X_{s}|Y}$ for all $s$. Therefore, $p_{Q|Y}$ satisfies the constraints in Eq. (30), so $I(Q;Y)\leq V_{1}$. Combining with Eq. (32) implies
@@
We now prove a slightly generalized version of Theorem 2. We show that the solution to the following optimization problem is non-decreasing and concave in $R$:
The slight generalization comes from replacing $\max$ in Eq. (12) by $\sup$, so that the result also holds for systems with infinite-dimensional sources where the supremum is not guaranteed to be achieved.
@@
To show concavity, consider any two points on the RB curve as defined by Eq. (34): $(R,I_{\text{RB}}(R))$ and $(R^{\prime},I_{\text{RB}}(R^{\prime}))$. For any $\epsilon>0$, there exist $Q$ and $Q^{\prime}$ such that
$I(Q;S|Y)$ $\leq R$ $I(Q;Y|S)$ $\geq I_{\text{RB}}(R)-\epsilon$
$I(Q^{\prime};S|Y)$ $\leq R^{\prime}$ $I(Q^{\prime};Y|S)$ $\geq I_{\text{RB}}(R^{\prime})-\epsilon$
Without loss of generality, suppose that both variables have the same set of outcomes $\mathcal{Q}$. Then, we define a new random variable $Q_{\lambda}$ with outcomes $\mathcal{Q}_{\lambda}=\{1,2\}\times\mathcal{Q}$, as well as a family of conditional distributions parameterized by $\lambda\in[0,1]$:
@@
In this way, we define $Q_{\lambda}$ via a disjoint convex mixture of $Q$ and $Q^{\prime}$ onto non-overlapping subspaces, with $\lambda$ being the mixing parameter. With a bit of algebra, it can be verified that, for every $\lambda$,
@@
$\leq\lambda R+(1-\lambda)R^{\prime}$
@@
$\geq\lambda I_{\text{RB}}(R)+(1-\lambda)I_{\text{RB}}(R^{\prime})-\epsilon$
Since $I_{\text{RB}}$ is defined via a maximization, we have
$I_{\text{RB}}(\lambda R+(1-\lambda)R^{\prime})\geq I(Q_{\lambda};Y|S)\\ \geq\lambda I_{\text{RB}}(R)+(1-\lambda)I_{\text{RB}}(R^{\prime})-\epsilon.$
Taking the limit $\epsilon\to 0$ proves the concavity. ∎